OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n3 + 2n is divisible by 3 for every positive integer n.

OpenStudy (anonymous):

have you ever done a proof by induction?

OpenStudy (solomonzelman):

Note that: \(\color{#000000 }{ \displaystyle n^3+2n=n(n^2+2) }\) we will split this into 3 categories: \(\color{#000000 }{ \displaystyle n\in 3k }\) \(\color{#000000 }{ \displaystyle n\in 3k+1 }\) \(\color{#000000 }{ \displaystyle n\in 3k+2 }\) Case 1: \(\color{#000000 }{ \displaystyle n^3+2n=n(n^2+2) \quad \Longrightarrow \quad (3k)\cdot [(3k)^2+2)] }\) (product of integers multiplied times 3, so that should be obvious) Case 2: \(\color{#000000 }{ \displaystyle n^3+2n \quad \Longrightarrow \quad (3k+1)^3+2(3k+1)}\) (expand and keep in mind that k is a natural number) Case 3: \(\color{#000000 }{ \displaystyle n^3+2n \quad \Longrightarrow \quad (3k+2)^3+2(3k+2)}\) (expand and keep in mind that k is a natural number) For cases 2 and 3, you are just expanding, and you will get an expression that will be divisible by 3 for any positive integer k.